OpenStudy (anonymous):

Rewrite the expression in simplest radical form. Simplify completely. 169^1/2

OpenStudy (anonymous):

169^1/2=\[\sqrt{169}\]

OpenStudy (jdoe0001):

\(\bf \large a^{\frac{n}{m}} = \sqrt[m]{a^n}\)

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

this one is very confusing :( :512^1/3

OpenStudy (jdoe0001):

same thing

OpenStudy (anonymous):

so it would be √512?

OpenStudy (jdoe0001):

\(\bf \huge a^{\frac{n}{m}} = \sqrt[m]{a^n}\)

OpenStudy (anonymous):

Not exactly. You would take the cube root in that case.

OpenStudy (anonymous):

a^n/m

OpenStudy (anonymous):

3√512 like this

OpenStudy (anonymous):

1=n 3=m Yes!

OpenStudy (anonymous):

\[\sqrt[3]{512}\]

OpenStudy (anonymous):

^Technically it should look like that, but I think you get the general idea.

OpenStudy (anonymous):

ok how would i type that on my keyboard

OpenStudy (jdoe0001):

\(\bf \huge \sqrt[\color{red}{3}]{512}\)

OpenStudy (anonymous):

^Thanks for making it bigger.